OpenStudy (anonymous):

if h(x)=x^2-5, find h(-3) a.-14 b.15 c.-11 d.4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

put -3 to place where you see x in the formula (-3)^2-5=4 (D)

jhonyy9 (jhonyy9):

9-5=4

jhonyy9 (jhonyy9):

d. is right sure

OpenStudy (anonymous):

d